OpenStudy (el_arrow):

find a formula for the general term of the sequence (-4,9/2,-16/3,25/4,-36/5,...)

OpenStudy (el_arrow):

i tried doing (n^2)/n+1* (-1)^n but i am not sure

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} &-4,\dfrac{9}{2},\dfrac{-16}{3},\dfrac{25}{4},\dfrac{-36}{5}\hspace{.33em}\\~\\ \implies& \dfrac{(-1)^{2+1}(2)^2}{2-1},\dfrac{(-1)^{3+1}(3)^2}{3-1},\\~\\&\dfrac{(-1)^{4+1}(4)^2}{4-1},\dfrac{(-1)^{5+1}(5)^2}{5-1},\dfrac{(-1)^{6+1}(6)^2}{6-1}\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (el_arrow):

so the exponent for -1 is n+1

OpenStudy (el_arrow):

i got confused because i thought that every sequence began with 1 but i guess not

OpenStudy (mathmath333):

ur formula is this ? \(\large \color{black}{\begin{align} \dfrac{(n^2)}{n+1}\cdot (-1)^n\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (el_arrow):

yeah thats the one i was using

OpenStudy (mathmath333):

lets work first on numerators only think a formula only for this sequence first, \(\large \color{black}{\begin{align} \quad-4,\quad 9,\quad-16,\quad36\cdots \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (el_arrow):

you forgot the 25

OpenStudy (el_arrow):

it would be n^2 * (-1)^n+1

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} 4,9,16,25,36\cdots \hspace{.33em}\\~\\ \implies (n+1)^2 \hspace{.33em}\\~\\ -4,9,-16,25,-36\cdots \hspace{.33em}\\~\\ \implies ? \end{align}}\)

OpenStudy (el_arrow):

(n-1)^2

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} -4,9,-16,25,-36\cdots \hspace{.33em}\\~\\ \Large \implies (-1)^n(n+1)^2 \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

see if it is coorect for numerators

OpenStudy (el_arrow):

yeah its correct

OpenStudy (mathmath333):

ok now work for only denominators \(\large \color{black}{\begin{align} \quad 1,\quad2,\quad3,\quad4,\quad5\cdots \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (el_arrow):

it just n

OpenStudy (mathmath333):

yes now combine numerator and denominator

OpenStudy (el_arrow):

an =(-1)^(n)* ((n+1)^(2)/(n)

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \dfrac{(-1)^n(n+1)^2}{n} \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (el_arrow):

yes how do you write it neat like that?

OpenStudy (freckles):

put \frac{ }{ } inside \[ \]

OpenStudy (freckles):

then put stuff in the braces

OpenStudy (freckles):

\frac{(-1)^n \cdot (n+1)^2}{n} inside \[ \]

OpenStudy (mathmath333):

go to latex subject and their is a tutorial on latex by thomaster u can learn a lot from there

OpenStudy (freckles):

oops those thingy disappear i forgot :p

OpenStudy (mathmath333):

or u can use equation editor

OpenStudy (freckles):



OpenStudy (el_arrow):

\frac{-1^n/n}

OpenStudy (mathmath333):

>.\(\large \color{black}{\begin{align} \dfrac{(-1)^n(n+1)^2}{n} \hspace{.33em}\\~\\ \end{align}}\.)<

OpenStudy (mathmath333):

copy betweem >. and .<

OpenStudy (freckles):

hey mathmath333 do you ever have the problem of copying things on OpenStudy?

OpenStudy (el_arrow):

>.\(\large \color{black}{\begin{align} \dfrac{(-1)^n(n+1)^2}{n} \hspace{.33em}\\~\\ \end{align}}\.)<

OpenStudy (el_arrow):

lol i'll just draw them things

OpenStudy (freckles):

I think he wanted you to copy everything in between the >. and .< and not including those thingy though

OpenStudy (mathmath333):

remove the \(.\) before \(.)\)

OpenStudy (el_arrow):

\(\large \color{black}{\begin{align} \dfrac{(-1)^n(n+1)^2}{n} \hspace{.33em}\\~\\ \end{align}}\

OpenStudy (mathmath333):

it will ne better if u just go latex section and read the tutorial latex on there u will come back as a hero

OpenStudy (mathmath333):

u forget to apply \()\) at last

OpenStudy (el_arrow):

whats that under

OpenStudy (mathmath333):

find more subjects

OpenStudy (freckles):

thats a link to the group

OpenStudy (mathmath333):

^^

OpenStudy (el_arrow):

oh thanks freckles